Compute the present value of $600 paid in three years using the following discount rates: 6 percent in the first year, 7 percent in the second year, and 8 percent in the third year.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

600=P*(1.06)*(1.07)*(1.08)

600=P*1.224936

P=600/1.224936

=$489.82(Approx).